A hockey player is standing on his skates on a frozen pond when an opposing player, moving with a uniform speed of 5.0 m/s, skates by with the puck. After 1.90 s, the first player makes up his mind to chase his opponent. If he accelerates uniformly at 0.11 m/s2, determine each of the following. (a) How long does it take him to catch his opponent? (Assume the player with the puck remains in motion at constant speed.) (b) How far has he traveled in that time? Step 1 (a) We choose x = 0 to be the initial location of the first player, and t = 0 to be the instant when this player starts to chase his opponent. At this time, the opponent is (1.90 s)·(5.0 m/s) = 9.5 m in front of the first player. We are given that the first player accelerates uniformly at 0.11 m/s2 At time t > 0 the displacements of the first player and his opponent from the origin are xplayer = (x0)player + (v0)player t + 1 2 aplayer t2 = 0 + 0 + 1 2 m/s2 t2 and xopponent = (x0)opponent + (v0)opponent t + 1 2 aopponent t2 = m + m/s t + 0
Answer #1

Let first player catches his opponent after t seconds from his starting time.

Distance moved by first player in t seconds = (1/2)at2 = 0.055 t2   ..............................(1)

where a is acceleration = 0.11 m/s2

When the first player is about to start, his opponent is ahead of him by distance = 1.95 = 9.5 m

if the opponent moves with constant speed 5 m/s in time duration t seconds,

then the distance moved by opponent = 5t m

Hence the total distance moved by the first player to catch his opponent = ( 9.5 + 5 t ) m ..............(2)

By equating eqns.(1) and (2), we get the following quadratic equation,

0.055 t2 = 9.5 + 5 t .....................(3)

t2 = 173 + 91 t       or    t2 -91 t - 173 = 0 ........................(4)

solution of the above quadratic eqn.(4) is given by,

------------------------------------------------------------

(b) distance S moved by the first player is obtained from eqn.(1) as

S = 0.055t2 = 0.0559393 476 m